Circular No. 6664

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970508 D. S. Shepherd, M. R. Metzger, and S. R. Kulkarni, California Institute

of Technology, write: “We have monitored the region of GRB 970508 at 86.8 GHz with the Owens Valley Radio Observatory Interferometer (lowresolution configuration; synthesized beam of diameter ∼ 5′′) from May 10 to 15. We first mosaicked a 40′′ × 40′′ region centered on the x-ray source SAX J0653.8+7916 (IAUC 6656), including the optical counterpart (IAUC 6654, 6655), from observations made during May 10.92–11.38 UT (map rms 3.6 mJy/beam) and May 11.81–12.53 (map rms 2.6 mJy/beam); no obvious sources were detected in either map. We then observed a single field centered on α = 6^{h}53^{m}47^{s}.11, δ = +79^{o}16′08′′.2 (equinox 2000.0) during May 13.01–13.9 (map rms 1.3 mJy/beam) and May 15.22–15.90 (map rms 2.27 mJy/beam). The optical transient was within 15′′ of this pointing center, and thus there was no significant loss in the primary beam response; no source was detected in either observation; the 3σ upper limits are 3.9 and 6.8 mJy/beam, respectively.”

COMET C/1997 J2 (MEUNIER-DUPOUY) Improved parabolic orbital elements from MPEC 1997-K05:

T = 1998 Mar. 10.744 TT  ω = 122.^{o}579
Ω = 148.874 }2000.0
q = 3.05760 AU  i = 91.307
1997 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
May 12 10^{h}51.^{m}2 +73^{o}42.^{′}5 4.278 4.204 79.^{o}0 13.^{o}6 12.9
22 10 47.5 +73 09.4 4.301 4.142 74.2 13.6 12.8
June 1 10 48.6 +72 29.7 4.316 4.082 70.0 13.5 12.8
11 10 54.1 +71 46.2 4.320 4.022 66.3 13.4 12.7
21 11 03.3 +71 00.7 4.312 3.963 63.4 13.3 12.7
July 1 11 15.9 +70 14.2 4.291 3.905 61.2 13.2 12.6
11 11 31.48 +69 26.9 4.257 3.848 59.8 13.2 12.5
21 11 49.63 +68 38.6 4.211 3.793 59.2 13.3 12.4
31 12 10.15 +67 48.7 4.153 3.738 59.3 13.5 12.3
Aug. 10 12 32.86 +66 56.3 4.084 3.685 60.1 13.8 12.2
20 12 57.56 +66 00.1 4.007 3.634 61.5 14.2 12.1
30 13 24.09 +64 58.7 3.923 3.584 63.2 14.6 12.0
Sept. 9 13 52.24 +63 50.3 3.835 3.535 65.3 15.0 11.9
1997 May 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green